A candy store owner wishes to mix candy that sells for $3 per pound with candy selling for $4 per pound to make a mixture to sell for $3.60 per pound. How many pounds of each kind of cany should be used to make an 80 pound mixture?
Thank you

Answer by ptaylor(1894)

Let x=amount selling for $3 per lb
Then 80-x=amount selling for $4 per lb
Now we know that the total amount of dollars generated by selling the candy separately would have to equal the total amount of dollars generated by selling the 80 lb mixture. Therefore, our equation to solve is:
$3x+(80-x)($4)=(80)($3.60) simplifying, we get:
3x+320-4x=288
-x=-32 or
x=32 lb of $3 candy
80-x=80-32=48 lb of $4 candy

Hope this helps -----ptaylor